8. Using the definition of the Laplace transform compute . {3}(s).
Answer #1

defn of Laplate freeform 12 f(t)} st fuldt dell st L 3 Sesx IP 8x = 3 -st e .st 3 e 5 E = e ēslo) + S = 310**] riē no oo 3 12